OpenStudy (anonymous):

Tell whether the sequence is arithmetic. If it is, what is the common difference? 15. 1, 1 1/6, 1 1/3, 1 1/2

OpenStudy (anonymous):

yes -1/6, yes 1/6, yes 6, no (These are the options

OpenStudy (anonymous):

Yes, you add 1/6 to each preceding term. Hence, 1/6 is the common difference.

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

@mangnagelytis Can you help me with another?

OpenStudy (anonymous):

Sure, what is it?

OpenStudy (anonymous):

16. Bamboo plants grow rapidly. A bamboo plant is 130 inches tall. Tomorrow it will be 143 inches tall, the next day it will be 156 inches tall, and on the next day it will be 169 inches tall. Write a rule to represent the height of the bamboo plant as an arithmetic sequence. How tall will the plant be in 13 days? (1 point) A(n) = 130 + (n – 1)13; 286 inches A(n) = 130 + (n – 1)13; 299 inches A(n) = 130 + 13n; 286 inches A(n) = 130 + 13n; 299 inches

OpenStudy (anonymous):

143-130=13, so the common difference would be 13 in/day. The equation would be 130 (the initial height)+(n-1)13, substitute 13 (days) into the equation for n and you get 130+(12)13=130+156=286 in

OpenStudy (anonymous):

Okay, that makes sense! Thank you!

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

Mind if I ask you another? This is the last one. I promise.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

17. Crystal earns $5.50 per hour mowing lawns. a. Write a rule to describe how the amount of money m earned is a function of the number of hours h spent mowing lawns. b. How much does Crystal earn if she works 3 hours and 45 minutes? (1 point)

OpenStudy (anonymous):

I think it's A but I'm not sure.

OpenStudy (anonymous):

She earns $5.50 an hour, so the equation would be m(h)=5.5h. 45 minutes is 3/4, or 0.75 of an hour, so substitute 3.75 for h and you get 20.625, or $20.63.

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

You're welcome.